A bus leaves a station at 1 p.m., traveling west at an average rate of 44 mi/h. One hour later a second bus leaves the same station, traveling east at a rate of 48 mi/h. At what time will the two buses be 274 mi apart?
Answer by checkley75(3666)

44t+48(t-1)=274
44t+48t-48=274
92t=274+48
92t=322
t=322/92
t=3.5 hours after the first bus leaves.
1+3.5=4.5 hours or 4:30 the buses will be 274 miles apart.
proof
44*3.5+48*2.5=274
154+120=274
274=274